# Test-Time Degradation Adaption for Open-Set Image Restoration

### v3

@@ 4.3 Analysis Experiments
Figure 6: The visualizations w.r.t TDA, where the Referred/Guided Images are translated to Degraded Referred/Guided Images by TDA, whose domains align well with that of Test Degraded Images.
The effectiveness of TDA. Since PDM cannot achieve image dehazing without TDA, we show its effectiveness by observing whether it aligns the domain of generated images to that of the degraded image after the optimization during the denoising process. The results are presented in Fig.6, from which one could observe that TDA aligns the domain of generated images well with that of the degraded images. For example, the haze in the top row is uneven distributed in the test degraded image, and the same effect exists in the degraded referred/guided image. Besides, in the bottom row, the degree of darkness in the degraded referred/guided image is similar to that in the test degraded image. Overall, our TDA exhibits significant capability of domain alignment to the unknown and unseen degradations.
@@
Table 4: Ablation studies on AIR, where U.G. denotes using the same guidance loss of mean squared error throughout the denoising process. -F.S., -S.S., and -T.S. receptively denote to remove the First Stage, Second Stage, and Third Stage in AIR, using only the above U.G.. The results demonstrate the indispensable roles of the three stages for image dehazing.
Metrics U.G. -F.S. -S.S. -T.G. TDG
PSNR $\uparrow$ 20.42 20.22 20.43 21.13 21.38
SSIM $\uparrow$ 0.851 0.857 0.837 0.857 0.859
The effectiveness of AIR. Here, we first demonstrate the rationality of TDG in Fig.7, which exhibits a typical denoising process of the PDM, from which one could see the significant temporal dynamics, i.e., the images are generated from the unrecognizable contents to the perceptually rich contents, and finally the imperceptible details. These dynamics also imply the varying of the contents should be focused and restored in the test degraded image, and thus applying different guidance strategies could facilitate the image restoration. Besides, we present the results of ablation studies on AIR in Tab.4, which demonstrate the effectiveness of our designs on facilitating the image dehazing.
@@ 5 Conclusion
This paper explicitly explores the challenges posed by open-set scenarios, and formally defines the problem of open-set image restoration. To solve this problem, this work reveals its essence from the perspective of distribution shifts, and discovers the methodology of test-time adaption is adept at addressing this inherent disparities. Motivated by this, this work presented a test-time degradation adaptation framework for open-set image restoration, which is ingenious in following ways. First, It considers a pre-trained image diffusion model as the general pretraining model for solving various tasks of image restoration. Then, it introduces an adapter optimized during the test phase for adapting the pre-trained model to the unknown and unseen test degradations. Finally, it dynamically adjusts the guidance strategies following the denoising process to obtain a better restoration result. Through experiments on multiple degradations, this work demonstrate the effectiveness of the proposed method.
Potential Broader Impact

### v4

@@ 4.3 Analysis Experiments
The effectiveness of TDA. Since PDM cannot achieve image dehazing without TDA, we demonstrate its effectiveness by i) observing whether it aligns the domain of generated images to that of the degraded image after the optimization during the denoising process, and ii) comparing with different strategies of adversarial training for domain alignment. The visual results are shown in Fig.6, from which one could observe that TDA exhibits a significant capability of domain alignment to the unknown and unforeseen degradations. For example, the haze in the top row is uneven distributed in the test degraded image, and the same effects exist in the degraded auxiliary/guided image. The degree of darkness in the degraded auxiliary/guided image is similar to that in the test degraded image in the second row. In addition, we conduct experiments on different strategies of adversarial training implemented by BasicSR (Wang et al., 2022b), and present the results in Tab. 4. From the table, one could see that, except for WGAN, all strategies achieve the significant performance gains over the haze images. Namely, our adapter generalizes well to different strategies for domain alignment to the unknown and unseen degradations.
@@
Table 5: Ablation studies on guidance strategies, where U.G. denotes using the same guidance loss of mean squared error throughout the denoising process. -F.S., -S.S., and -T.S. denote to remove the First Stage, Second Stage, and Third Stage, receptively. The results demonstrate the indispensable roles of the three stages for image dehazing.
Metrics U.G. -F.S. -S.S. -T.S. AIR
PSNR $\uparrow$ 20.49 21.43 21.06 21.37 22.29
SSIM $\uparrow$ 0.869 0.873 0.865 0.871 0.877
The effectiveness of AIR. Here, we first demonstrate the rationality of AIR in Fig.7 which shows a typical denoising process of PDM. From the figure, one could see the significant temporal dynamics, i.e., the images are generated from the unrecognizable contents to the perceptually rich contents, and finally the imperceptible details. The dynamics also imply the varying of the contents should be focused and restored in the test degraded image, and thus applying different guidance strategies could facilitate image restoration. We present the ablation results in Tab. 5, which show the effectiveness of the three stages on facilitating image dehazing. In addition, we also conduct analysis experiments on stage divisions as shown in Tab. 6, from which one could see that both early and late dividing steps result in the inferior performances. In other words, a proper stage division consistent with the temporal dynamics could improve the performance of image restoration.
Table 6: Analysis experiments on divisions of the three stages. The left is the dividing step between the First and Second stages, and the right is the dividing step between the Second and Third stages. The results demonstrate a proper stage division could significantly improve the restoration.
Metrics 800 700 600 70 50 30
PSNR $\uparrow$ 21.89 22.29 21.11 21.76 22.29 22.01
SSIM $\uparrow$ 0.861 0.877 0.868 0.876 0.877 0.872
@@ 5 Conclusion
This paper explicitly explores the challenges posed by open-set scenarios, and formally defines the problem of open-set image restoration. To solve this problem, this work reveals its essence from the perspective of distribution shifts, and discovers the methodology of test-time adaptation is adept at addressing this inherent disparities. Motivated by this, we presented a test-time degradation adaptation framework for open-set image restoration, which is ingenious in the following ways. First, it considers a pre-trained image diffusion model as the general pretraining model for solving various tasks of image restoration. Second, it introduces an adapter optimized during the test phase for adapting the pre-trained model to the unknown and unseen test degradations. Third, it dynamically adjusts the guidance strategies following the denoising process to obtain better restoration results. Through experiments on multiple degradations, we demonstrate the effectiveness of our designs.
Impact Statement
@@
Acknowledgment
The authors would like to thank the anonymous reviewers and area chair for their valuable comments and constructive suggestions to improve the quality of this paper. This work was supported in part by NSFC under Grant 62176171, U21B2040; in part by the Fundamental Research Funds for the Central Universities under Grant CJ202303; and in part by Sichuan Science and Technology Planning Project under Grant 24NSFTD0130.